# Augment then Smooth: Reconciling Differential Privacy with Certified Robustness

### v1

@@ Abstract
Machine learning models are susceptible to a variety of attacks that can erode trust in their deployment. These threats include attacks against the privacy of training data and adversarial examples that jeopardize model accuracy. Differential privacy and randomized smoothing are effective defenses that provide certifiable guarantees for each of these threats, however, it is not well understood how implementing either defense impacts the other. In this work, we argue that it is possible to achieve both privacy guarantees and certified robustness simultaneously. We provide a framework called DP-CERT for integrating certified robustness through randomized smoothing into differentially private model training. For instance, compared to differentially private stochastic gradient descent on CIFAR10, DP-CERT leads to a 12-fold increase in certified accuracy and a 10-fold increase in the average certified radius at the expense of a drop in accuracy of 1.2%. Through in-depth per-sample metric analysis, we show that the certified radius correlates with the local Lipschitz constant and smoothness of the loss surface. This provides a new way to diagnose when private models will fail to be robust.
@@ 1 Introduction
Machine learning (ML) models are becoming increasingly trusted in critical settings despite an incomplete understanding of their properties. This raises questions about the trustworthiness of those models, encompassing aspects such as privacy, robustness, and more. Society at large might expect all of these properties to hold simultaneously as ML’s influence on everyday life expands, but each aspect is challenging enough that scientists and practitioners still mostly grapple with them individually. Relatively little research has been done on the intersectionality of trustworthy ML requirements, since each aspect seems to push us in orthogonal research directions.
@@
We aim to reconcile two key objectives of trustworthy ML, namely privacy and robustness. Privacy in the context of ML manifests as the requirement that a model does not leak information about the data it was trained on Papernot et al. (2018), such as revealing whether or not certain data points were included in the training dataset Shokri et al. (2017) or what characteristics they exhibit Fredrikson et al. (2015). In our study, robustness refers to the requirement that a model’s prediction should not change when its test inputs are perturbed, even in the worst case when perturbations are chosen adversarially Biggio et al. (2013); Szegedy et al. (2013); Goodfellow et al. (2014).
@@
The current gold standard for providing privacy guarantees is differential privacy (DP) Dwork and Roth (2014). In ML, DP produces mathematically rigorous privacy guarantees by limiting the impact of each individual training data point on the final model. This is achieved by clipping per-sample gradients, and adding a well-calibrated amount of noise to all model updates. Clipping serves to bound the sensitivity of the training algorithm, while the addition of noise ensures that training will be more likely to output similar models whether any of the individual data points are added to or removed from the training dataset. However, clipping and adding noise can impede the convergence of models Tramèr and Boneh (2021) and yield decision boundaries that are less smooth Hayes et al. (2022), negatively impacting robustness Fawzi et al. (2018).
These findings call for integrating robustness measures into private training, yet this remains challenging because most methods to increase robustness use random or adversarial augmentations of training data points, which both conceptually and practically do not align well with DP training. Conceptually, augmenting an input increases the sensitivity of private training to it, and thereby provides additional avenues for information leakage. From a practical viewpoint, since gradients are computed on a per-example basis for DP, augmentations drastically increase the time and memory costs of training.
@@
To bridge the gap between robust and private ML model training, we evaluate the certified robustness of private models and improve it by integrating state-of-the-art techniques Wong and Kolter (2018); Raghunathan et al. (2018) with DP training. CR provides probabilistic guarantees that perturbations of a certain magnitude will not change a model’s prediction, regardless of what attack strategy (known or yet unknown) is used to modify the test inputs, and thereby provides future-proof robustness guarantees. A common approach for certifying robustness is randomized smoothing, where a classifier’s outputs are averaged over a distribution surrounding the test point Lecuyer et al. (2019); Li et al. (2019); Cohen et al. (2019).
While DP and CR are the most promising standards for providing future-proof privacy and robustness guarantees respectively, their intersection has seen little attention. Recent works Phan et al. (2019, 2020); Tang et al. (2022) propose adding noise or adversarial examples while training to improve CR guarantees, but lack the flexibility to incorporate state-of-the-art methods for non-private training Wong and Kolter (2018); Raghunathan et al. (2018), and usually rely on training additional network components. We aim to provide CR guarantees within the standard DP training framework, overcoming several challenges in doing so.
@@
We study the possible pitfalls of combining DP and CR in a systematic manner. Through our analysis and ablation studies combining randomized smoothing techniques with DP training, we show that standard DP training of ML models is insufficient to provide strong CR results. We propose DP-CERT, an adaptable framework for integrating CR into standard DP training which effectively incorporates augmentations while managing the additional privacy risks. Compared to private training without augmentations, DP-CERT achieves better robustness on MNIST, Fashion-MNIST, and CIFAR10, and even surpasses the state-of-the-art for robustness on the latter dataset under the same privacy guarantee. Finally, we analyze CR on a per data point basis rather than averaged across test datasets. Using the gradient norm, Hessian spectral norm, and local Lipschitz constant, we find that the certifiable radius has a negative log-linear correlation with these quantities, and compare their distributions across training methods. We conclude with concrete recommendations of best practices for the community to achieve CR and DP simultaneously.
@@ 2 Preliminaries
Consider a classification task with $Y$ classes from a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{1,{...},Y\}$ denote the $i$-th input and label. Let $f_{\theta}:\mathbb{R}^{d}\xrightarrow{}\{1,{...},Y\}$ be a neural network with parameters $\theta$, and $F_{\theta}$ denote the soft classifier which outputs the probability distribution, such that $f_{\theta}(x)=\operatorname*{arg\,max}_{y\in\{1,{...},Y\}}F_{\theta}(x)_{y}$, where $F_{\theta}(x)_{y}$ denotes the model probability of $x$ being a member of class $y$.
@@
We rely on the rigorous framework of differential privacy (DP) Dwork et al. (2006) to obtain models with privacy guarantees. DP ensures that a model’s weights at the end of training will be similar in distribution whether or not a particular data point was included in the training set. More formally, let $D$ and $D^{\prime}$ be two potential training datasets for a model $f_{\theta}$ that differ in only one data point. The training mechanism $M$ guarantees $(\varepsilon,\delta)$-DP if for all possible sets of outcomes $S$ of the training process, it holds that $\mathrm{Pr}\left[M(D)\in S\right]\leq e^{\varepsilon}\mathrm{Pr}\left[M(D^{\prime})\in S\right]+\delta$. The parameter $\varepsilon$ specifies the privacy level, with smaller $\varepsilon$ yielding higher privacy, while $\delta$ quantifies the probability of the algorithm violating the $\varepsilon$ privacy guarantee.
@@
To obtain a differentially private variant of stochastic gradient descent (SGD), two modifications need to be made Abadi et al. (2016). First, the individual gradients of each data point are clipped to a norm $C$ to limit the sensitivity of the model update caused by each data point. Second, choosing a noise level $\rho$, noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$ is added to the aggregated gradients to prevent the changes to the model from revealing too much information about individual data points. We detail the resulting algorithm DPSGD (1) and a more thorough introduction to DP in Appendix A.1.
Certified Robustness.
Adversarial examples are a well-studied phenomenon in ML, in which an input to a model is perturbed in ways that do not alter its semantics yet cause the model to misclassify the perturbed input Biggio et al. (2013); Szegedy et al. (2013); Goodfellow et al. (2014). Formally, for a given labeled datapoint $(x,y)$ and classifier $f$, an $(\ell_{p},\zeta)$-adversary aims to create an adversarial example $x^{\prime}$ such that $\|x^{\prime}-x\|_{p}<\zeta$ and $f(x^{\prime})\neq y$. Despite much research, the most common defense against adversarial examples remains adversarial training Goodfellow et al. (2014); Zhang et al. (2019b); Rebuffi et al. (2021). While adversarial training improves robustness to known algorithms for finding adversarial examples, it does not guarantee that a model will be robust to all adversarial examples (e.g., those crafted with other attack algorithms). This motivates the development of techniques that can provide certifiable guarantees of robustness to adversarial examples by providing a lower bound $r$ on the distance between a correctly classified input and any adversarial example that may be misclassified Wong and Kolter (2018); Raghunathan et al. (2018). This lower bound is also known as the certification radius.
Randomized Smoothing.
One popular approach for establishing certified robustness (CR) guarantees is through probabilistic robustness verification which, with high probability, verifies that no adversarial examples exist within a certain radius of the original input Li et al. (2023). The most commonly studied method for providing a probabilistic robustness verification is through smoothing a classifier Lecuyer et al. (2019); Li et al. (2019); Cohen et al. (2019) by averaging the class predictions of $f$ using a smoothing distribution $\mu$,
[…]
where $\mathbb{I}[a,b]=1\iff a=b$ and 00 otherwise Li et al. (2023). As computing the integral in Equation (1) is intractable, Monte Carlo sampling is used. We denote the approximation of $\hat{g}$ given by Monte Carlo sampling as $g$. One can certify at different radii through the choice of smoothing distribution $\mu$. Smoothed classifiers are evaluated in terms of their certified accuracy—the fraction of samples correctly classified when certifying robustness at a given radius $r$.
A tight $\ell_{2}$ radius was obtained by Cohen et al. (2019) when using isotropic Gaussian noise $\mu=\mathcal{N}(x,\sigma^{2}\mathbf{I})$, where $\sigma$ is a hyperparameter that controls a robustness/accuracy tradeoff. In particular, Cohen et al. (2019) proved that for any base classifier $f$, the Gaussian smoothed classifier $g$ is robust around an input $x$ with radius $r=\frac{\sigma}{2}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B}))$ where $p_{A}$ and $p_{B}$ denote the probabilities of $c_{A}$ and $c_{B}$, the most and second-most probable classes returned by $g(x)$, and $\Phi^{-1}$ is the inverse of the standard Gaussian CDF. In fact the exact probabilities $p_{A}$ and $p_{B}$ are not needed and one can use lower $\underline{p_{A}}\leq p_{A}$ and upper $\overline{p_{B}}\geq p_{B}$ bounds instead, approximated by Monte Carlo sampling.
The output of the smoothed classifier $g(x)$ is approximated by aggregating the predictions of a base classifier $f(x+\eta)$ for $\eta\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. As a high dimensional standard Gaussian assigns almost no mass near its mean 00, ensuring that $g(x)$ is accurate at large certification radii requires the base classifier $f$ to be accurate on Gaussian perturbed data Gao et al. (2022).
3 Method
Training machine learning models to be both differentially private and certifiably robust poses several challenges. The gradient clipping and noise addition used in DPSGD harms the convergence rate of training Chen et al. (2020); Tramèr and Boneh (2021); Bu et al. (2021), while restrictive privacy budgets may further require stopping training prior to convergence. Robustness on the other hand suffers for models that are not converged, as having large gradients at test points makes finding adversarial examples easier Fawzi et al. (2018).
Another challenge surfaces around the use of adversarial training Goodfellow et al. (2014) or augmentations of datapoints Cohen et al. (2019) along with DPSGD. As shown by Cohen et al. (2019), data augmentations used for training can enhance a model’s CR, however, it is crucial to ensure that augmented data points do not leak private information about the original. Previous works on the combination of DP and CR have proposed adding noise or adversarial examples during training, but deviate from the standard DPSGD template to address the privacy risks Phan et al. (2019, 2020); Tang et al. (2022). These approaches add trainable model components increasing the overall complexity Phan et al. (2019); Tang et al. (2022), or lack the flexibility to incorporate the latest advancements in adversarial training methods Phan et al. (2020). For a more detailed description of these related works and comparison to our method, please see Appendix B.
We aim to make CR feasible within the standard training procedure of DPSGD, with state-of-the-art convergence and proper accounting for additional privacy risks by introducing the DP-CERT framework. In this section, we describe DP-CERT, how it effectively manages training with augmented samples while preserving privacy, and how it enables the integration of recent advancements in adversarial training and regularizers to enhance certifiable robustness Salman et al. (2019); Li et al. (2019); Zhai et al. (2020). Our training framework consists of three stages, summarized in Figure 1: augmentation multiplicity as the foundational stage, plus regularization and adversarial training as two optional stages. After the model is trained, randomized smoothing is used at inference time. We present four instantiations of the framework: DP-Gaussian, DP-SmoothAdv, DP-Stability, and DP-MACER, employing different techniques at each stage.
Augmentation Multiplicity.
For each data point $(x_{i},y_{i})$, we obtain $K$ augmented data points $(x_{i}^{j},y_{i})$, where $j\in\{1,...,K\}$ and $x_{i}^{j}$ is the $j$-th augmented data point. For notational convenience, we use $x_{i}^{0}$ to denote the original data point $x_{i}$. As shown by Cohen et al. (2019), training with Gaussian data augmentation can enhance a model’s certified robustness. When not using adversarial training, we define $x_{i}^{j}=x_{i}+\eta_{j}$, $\eta_{j}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ for $j\neq 0$.
An important component of our DP-CERT is how we handle training with augmented data points. We adopt augmentation multiplicity, introduced in De et al. (2022) and previously unused in studies of CR for DP, which involves averaging the gradients of multiple augmentations of the same training sample before clipping. Since all downstream impact to the model weights from sample $x_{i}$ is contained in this averaged gradient, clipping it provides a finite sensitivity as required for the Sampled Gaussian Mechanism used in DPSGD Mironov et al. (2019), and no additional privacy cost is incurred. The model updates can be expressed as follows
$\theta^{t+1}=\theta^{t}-\lambda_{t}\Biggl{[}\frac{1}{B}\sum_{i\in B_{t}}\textrm{clip}_{C}\Biggl{(}\frac{1}{K+1}\sum_{j=0}^{K}\nabla_{\theta^{t}}L_{\text{CE}}(x_{i}^{j},y_{i})\Biggr{)}+\frac{\rho C}{B}\xi\Biggr{]}\text{.}$ (2)
$\theta^{t}$ denotes the model parameters at iteration $t$, $\lambda_{t}$ is the learning rate, $B$ is the batch size, $C$ is the clipping bound, $K$ is the number of augmentations, $\rho$ is the noise multiplier, $\xi\sim\mathcal{N}(0,\textbf{I})$, and $\nabla_{\theta^{t}}L_{\text{CE}}(x_{i}^{j},y_{i})$ is the gradient with respect to data point $(x_{i}^{j},y_{i})$. Note that $j$ starts from 0, which means we include the original samples along with the augmented ones in model training.
Regularization.
We propose adapting stability and consistency regularization to private training in order to minimize the distance between the output probability of the original and augmented examples, hereby improving the robustness to input noise. Stability training Li et al. (2019) adds a smoothed cross-entropy loss as regularization. Inspired by TRADES Zhang et al. (2019a), we instead use the Kullback–Leibler (KL) divergence with a hyperparameter $\gamma$ controlling the strength of the regularization as:
$L_{\text{stability}}(x_{i},y_{i})=\sum_{j}L_{\textrm{CE}}(x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}F_{\theta}(x_{i})||F_{\theta}(x_{i}^{j})\big{)}\text{.}$ (3)
Consistency regularization Jeong and Shin (2020) is a similar technique that instead minimizes the KL divergence between $\hat{F}_{\theta}(x_{i})$ and $F_{\theta}(x_{i})$, where $\hat{F}_{\theta}(x)=\frac{1}{K}\sum_{j}F_{\theta}(x_{i}^{j})$ is the average output probability of all smoothed samples. The loss can be expressed as
$L_{\text{consistency}}(x_{i},y_{i})=\sum_{j}L_{\textrm{CE}}(x_{i}^{j},y_{i})+\gamma D_{\text{KL}}\big{(}\hat{F}_{\theta}(x_{i})||F_{\theta}(x_{i}^{j})\big{)}\text{.}$ (4)
Additionally, we propose integrating MACER Zhai et al. (2020), an alternative training modification to directly optimize the certified accuracy at larger robustness radii without requiring the costly process of adversarial training. MACER achieves this by decomposing the error of a smoothed classifier into a classification error term and a robustness error term, the latter reflecting whether or not the smoothed classifier was able to certify robustness for a given radius.
Adversarial Training.
To achieve better certified accuracy, we incorporate adversarial training by deploying existing attacks to create adversarial examples. Specifically, we integrate SmoothAdv Salman et al. (2019) into private training, which, given original data $(x,y)$, optimizes
$\operatorname*{arg\,max}_{\|x^{\prime}-x\|_{2}\leq\epsilon}\Big{(}-\log\underset{\eta\sim N(0,\sigma^{2}I)}{\mathbb{E}}[F_{\theta}(x^{\prime}+\eta)_{y}]\Big{)},$ (5)
to find an $x^{\prime}$ $\epsilon$-close to $x$ that maximizes the cross entropy between $g_{\theta}(x^{\prime})$ and label $y$. Using Monte Carlo sampling, Objective (5) can be optimized by iteratively computing the approximate gradient
$\nabla_{x^{\prime}}\Big{(}-\log\Big{(}\frac{1}{K}\sum_{j=1}^{K}F_{\theta}(x^{\prime}+\eta_{j})_{y}\Big{)}\Big{)}.$ (6)
3.1 Metrics for Interpreting Robustness
To elicit some insights into why certain training methods may produce better-performing models than others, we investigate several per-data point metrics associated with robustness, the input gradient norm, input Hessian spectral norm, and local-Lipschitz constant, and study their relationships with CR. The first two metrics measure the local smoothness of the loss landscape with respect to the input space. Taylor’s approximation can be used to show a direct link between these two metrics and the worst-case change in loss from small input perturbations. Due to this connection, prior works directly regularized them in order to train more robust models Hoffman et al. (2019); Jakubovitz and Giryes (2018); Moosavi-Dezfooli et al. (2019).
Gradients and Hessians are highly local quantities that are only connected to robustness through Taylor’s approximation at small radii around the input data point. Consequently, they may not be informative at larger radii used to certify robustness. Thus, we also compare models using an empirical estimate of the average local Lipschitz constant of the model’s penultimate layer. By viewing the network as a feature extractor composed with a linear classifier, using the penultimate layer captures the worst-case sensitivity of the feature extractor to perturbations of the data. This metric was initially proposed by Yang et al. (2020) to investigate adversarial robustness and is given by
$\frac{1}{n}\sum_{i=1}^{n}\max_{x_{i}^{\prime}\in B_{\infty}(x,\zeta)}\frac{\|f(x_{i})-f(x_{i}^{\prime})\|_{1}}{\|x_{i}-x_{i}^{\prime}\|_{\infty}},$ (7)
where the maximum is approximated in the same manner as is used for adversarial example generation, typically projected gradient descent (PGD) Madry et al. (2018).
4 Conclusion
We achieve better certified robustness with DPSGD training through augmentations and randomized smoothing, reconciling two crucial objectives for trustworthy ML, namely privacy and robustness. To overcome the theoretical and practical challenges that arise from the combination of both approaches, we rely on state-of-the-art DP training with augmentations that does not incur additional privacy costs. We employ various regularizations, and adversarial training methods to enhance robustness. Our resulting DP-CERT framework is modular and supports multiple combinations of these methods. Through our extensive experimental study, we confirm that DPSGD training alone, even with state-of-the-art convergence, does not provide satisfactory certified robustness. However, introducing a small number of computationally inexpensive augmentations into training, such as adding Gaussian noise, suffices to yield strong privacy protection and certified robustness. By thoroughly analyzing per-sample metrics, we show that the certified radius correlates with the local Lipschitz constant and smoothness of the loss surface; this opens a new path to diagnosing when private models will fail to be robust. To conclude, our findings yield concrete recommendations for the community to simultaneously achieve CR and DP, providing a valuable contribution towards more trustworthy ML. When training from scratch, Gaussian augmentations (not adversarial) should be used with DPSGD, and randomized smoothing applied at inference time. For fine-tuning pretrained models, adding stability regularization also helps accuracy, and leads to much lower local Lipschitz constants.
Acknowledgments.
Appendix A Differential Privacy Background
A.1 Differential Privacy and DPSGD
Differential Privacy (DP) Dwork et al. (2006) is a formal framework that aims to provide mathematical guarantees on the privacy of individual data points in a dataset, while allowing one to learn properties over the entire dataset. More formally a randomized mechanism $M$ fulfills $(\varepsilon,\delta)$-DP if
where $D$ and $D^{\prime}$ are neighboring datasets (i.e. datasets differing in only one data point), $S$ is the set of output values of $M$, and $\epsilon$ and $\delta$ are privacy parameters.
$\varepsilon$ indicates the level of obtained privacy, with a smaller value of $\varepsilon$ indicating a stronger privacy guarantee. $\delta$ accounts for the probability that the algorithm can violate the privacy guarantee, i.e., not stay within the specified $\varepsilon$. A larger value of $\delta$ increases the chance of privacy leakage.
The most popular algorithm for implementing DP guarantees in ML is differentially private stochastic gradient descent (DPSGD) Abadi et al. (2016). It extends the standard SGD algorithm with two additional steps, namely gradient clipping and noise addition. While the former bounds the sensitivity of the model update, the latter implements the privacy guarantee by preventing the gradients from revealing too much information about individual data points. More formally, let $\theta_{t}$ denote the model parameters at training iteration $t$. At each iteration $t$, DPSGD computes the gradient of the loss function with respect to $\theta_{t}$ at an individual data point $x$ as $\nabla_{\theta}L(\theta_{t},x)$. The data point’s gradient is then clipped to a maximum norm $C$ using the operation $\text{clip}\left(\nabla_{\theta}L(\theta_{t},x),C\right)$, which replaces the gradient with a vector of the same direction but smaller magnitude if its norm exceeds $C$. The clipped gradient from all datapoints in a batch are aggregated, then perturbed by adding random noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$, where $\rho$ is the noise scale parameter. We detail DPSGD in Algorithm 1.
Algorithm 1 Standard DPSGD, adapted from Abadi et al. (2016).
@@ A.1 Differential Privacy and DPSGD
0:  Private training set $D=\{(x_{i},y_{i})\mid i\in[N_{\text{prv}}]\}$, loss function $L(\theta_{t},x)$, Parameters: learning rate $\eta_{t}$, noise scale $\rho$, group size $B$, gradient norm bound $C$.
@@
4:     For each $i\in B_{t}$, compute ${\mathbf{g}}_{t}(x_{i})\leftarrow\nabla_{\theta}L({\theta}_{t},x_{i})$Compute per-sample gradients}
@@
7:      ${\theta}_{t+1}\leftarrow{\theta}_{t}-\eta_{t}\tilde{{\mathbf{g}}}_{t}$Gradient Descent}

### v3

@@ Abstract
Machine learning models are susceptible to a variety of attacks that can erode trust, including attacks against the privacy of training data, and adversarial examples that jeopardize model accuracy. Differential privacy and certified robustness are effective frameworks for combating these two threats respectively, as they each provide future-proof guarantees. However, we show that standard differentially private model training is insufficient for providing strong certified robustness guarantees. Indeed, combining differential privacy and certified robustness in a single system is non-trivial, leading previous works to introduce complex training schemes that lack flexibility. In this work, we present DP-CERT, a simple and effective method that achieves both privacy and robustness guarantees simultaneously by integrating randomized smoothing into standard differentially private model training. Compared to the leading prior work, DP-CERT gives up to a 2.5$\times$ increase in certified accuracy for the same differential privacy guarantee on CIFAR10. Through in-depth per-sample metric analysis, we find that larger certifiable radii correlate with smaller local Lipschitz constants, and show that DP-CERT effectively reduces Lipschitz constants compared to other differentially private training methods. Code is available at github.com/layer6ai-labs/dp-cert.
@@ 1 Introduction
Machine learning (ML) models are becoming increasingly trusted in critical settings despite an incomplete understanding of their properties. This raises questions about the trustworthiness of those models, encompassing aspects such as privacy, robustness, and more. Society at large might expect all of these aspects to be accounted for simultaneously as ML’s influence on everyday life expands, but scientists and practitioners still mostly grapple with each aspect individually.
@@
We aim to reconcile two key objectives of trustworthy ML, namely privacy and robustness. Privacy in the context of ML manifests as the requirement that a model does not leak information about the data it was trained on (Papernot et al., 2018), such as revealing whether or not certain data points were included in the training dataset (Shokri et al., 2017) or what characteristics they exhibit (Fredrikson et al., 2015). In our study, robustness refers to the requirement that a model’s prediction should not change when its inputs are perturbed at test-time, even in the worst case of adversarially chosen perturbations.
@@
The current gold standard for providing privacy guarantees is differential privacy (DP) (Dwork & Roth, 2014). In ML, DP produces mathematically rigorous privacy guarantees by limiting the impact of each training data point on the final model. This is achieved by clipping per-sample gradients, and adding a well-calibrated amount of noise to all model updates. Clipping limits the sensitivity of the training algorithm, while the addition of noise ensures that training will be more likely to output similar models when any individual data point is added to or removed from the training dataset.
@@
To quantify robustness, we focus on test-time certified robustness (CR) (Wong & Kolter, 2018; Raghunathan et al., 2018), which provides probabilistic guarantees that perturbations of a certain magnitude will not change a model’s prediction, regardless of what attack strategy is used to modify the inputs. A common approach for certifying robustness is randomized smoothing, where a classifier’s outputs are averaged over a distribution surrounding the data point (Lecuyer et al., 2019; Li et al., 2019; Cohen et al., 2019).
Unfortunately, the two aims of providing DP and CR guarantees are in conflict, both empirically and conceptually. The clipping and noise addition used in DP training can impede the convergence of models (Tramèr & Boneh, 2021) and yield decision boundaries that are less smooth (Hayes et al., 2022), negatively impacting robustness (Fawzi et al., 2018). There is also substantial empirical evidence that providing privacy guarantees is in tension with robustness (Boenisch et al., 2021; Tursynbek et al., 2021). Integrating robustness measures into private training remains challenging because most methods to increase robustness use random or adversarial augmentations of training data points, which do not align well with DP training. Conceptually, augmenting an input increases the sensitivity of private training to it, and thereby provides additional avenues for information leakage. From a practical viewpoint, since gradients are computed on a per-example basis for DP, augmentations drastically increase the time and memory costs of training.
@@
We study the possible pitfalls of combining DP and CR. Through our analysis and ablation studies combining randomized smoothing techniques with DP training, we show that standard DP training of ML models is insufficient to provide strong CR results. Surprisingly, we find that prior complex methods used to regularize models to improve CR are unnecessary (Phan et al., 2019; 2020; Tang et al., 2022). Instead, we propose DP-CERT, a straightforward and adaptable framework for integrating CR into standard DP training which effectively incorporates augmentations while managing the additional privacy risks. Compared to the few existing approaches for integrating DP and CR, our proposed DP-CERT method is simpler, more flexible, and does not require extra network components. This decreased complexity is important for the widespread adoption of private and robust ML. DP-CERT even surpasses the state-of-the-art robustness guarantees on CIFAR10 under equivalent privacy guarantees.
To provide some insight into the mechanisms by which regularization methods improve CR for private models, we analyze CR on a per-data point basis. Using the gradient norm, Hessian spectral norm, and local Lipschitz constant, we find that the certifiable radius has a negative log-linear correlation with these quantities, and compare their distributions across methods. We conclude with concrete recommendations of best practices for the community to achieve CR and DP simultaneously.
@@ 2 Preliminaries
We consider a $Y$-class classification task with a dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\{1,{...},Y\}$. Let $f_{\theta}:\mathbb{R}^{d}\xrightarrow{}\{1,{...},Y\}$ be a neural network classifier with parameters $\theta$, and $F_{\theta}$ denote the soft classifier which outputs the probability distribution, such that $f_{\theta}(x)=\operatorname*{arg\,max}_{y\in\{1,{...},Y\}}F_{\theta}(x)_{y}$, where $F_{\theta}(x)_{y}$ denotes the model probability of $x$ being a member of class $y$.
@@
We rely on the rigorous framework of differential privacy (DP) (Dwork et al., 2006) to obtain models with privacy guarantees. DP ensures that a model’s weights at the end of training will be similar in distribution whether or not a particular data point was included in the training set. More formally, let $D$ and $D^{\prime}$ be two potential training datasets for a model $f_{\theta}$ that differ in only one data point. The training mechanism $M$ guarantees $(\varepsilon,\delta)$-DP if, for all possible sets of outcomes $S$ of the training process, it holds that $\mathrm{Pr}\left[M(D)\in S\right]\leq e^{\varepsilon}\mathrm{Pr}\left[M(D^{\prime})\in S\right]+\delta$. The parameter $\varepsilon$ specifies the privacy level, with smaller $\varepsilon$ yielding higher privacy, while $\delta$ quantifies the probability of a catastrophic failure of privacy (Kifer et al., 2022).
@@
To obtain a differentially private variant of stochastic gradient descent (SGD), two modifications need to be made (Abadi et al., 2016). First, the individual gradients of each data point are clipped to a norm $C$ to limit the sensitivity of the model update caused by each data point. Second, choosing a noise level $\rho$, noise from $\mathcal{N}(0,\rho^{2}C^{2}\mathbf{I})$ is added to the aggregated gradients to prevent the changes to the model from revealing too much information about individual data points. The resulting algorithm DPSGD is detailed in Algorithm 1. We provide a more thorough introduction to DP and explanation of Algorithm 1 in Appendix A.1.
Algorithm 1 Standard DPSGD, adapted from (Abadi et al., 2016).
@@
0:  Private training set $D=\{(x_{i},y_{i})\mid i\in[N_{\text{prv}}]\}$, loss function $L(\theta_{t};x,y)$, Parameters: learning rate $\lambda_{t}$, noise scale $\rho$, group size $B$, gradient norm bound $C$.
@@
4:     For each $i\in B_{t}$, compute ${\mathbf{g}}_{t}(x_{i})\leftarrow\nabla_{\theta}L({\theta}_{t};x_{i},y_{i})$Compute per-sample gradients}
@@
7:      ${\theta}_{t+1}\leftarrow{\theta}_{t}-\lambda_{t}\tilde{{\mathbf{g}}}_{t}$Gradient descent}